# Mitigating Negative Side Effects in Multi-Agent Systems Using Blame Assignment

### v1

@@ 5 Penalty Decomposition via Blame Assignment
Blame Estimation. Blame (credit) assignment estimates an agent’s contribution to a global penalty (or reward) using counterfactual states. The blame $B_{i}$ (Eqn. 1) is calculated for each agent $i$ in the joint state $s$, based on the difference between the current NSE penalty and the minimum NSE penalty that could have been achieved by the agent. The existing credit assignment techniques calculate counterfactual states by considering all state features. In our setting, NSE occurrence is determined only by the dynamic global features $(\vec{f}_{gd})$. Therefore, the counterfactuals must be calculated only over $\vec{f}_{gd}$. Generating counterfactual states for NSEs by considering all the state features could result in incorrect attribution. Consider multiple warehouse robots, with some carrying large shelves, navigating in a narrow corridor. The NSE of blocked paths for human access is determined by number of agents carrying large shelves. Generating counterfactual states traditionally will involve changing agent location but that does not provide information on NSE associated with multiple agents carrying large shelves. Hence, counterfactuals must be generated over NSE-relevant state features. We present a modified formulation of credit assignment that uses counterfactual neighbors instead of counterfactual states.
Definition 4.
The counterfactual neighbors of a joint state $s$, denoted by $s_{c}$, is the set of all states that vary only in the values of dynamic global features $(\vec{f}_{gd})$ while all other feature values are same as in $s$.
Counterfactual neighbors are therefore a subset of counterfactual states. Since the NSE penalty depends only on the state in our formulation, we calculate counterfactual neighbors over the state variables. The set of counterfactual neighbors that are reachable from the start state in the environment are referred to as valid counterfactual neighbors and are denoted by ${s}_{c}^{v}$.
Definition 5.
Agent-specific counterfactual neighbors, denoted by $s^{i}_{c}\!\subset s_{c}$, are states that differ in those dynamic global features that can be controlled by agent $i$, while other feature values are same as in the current joint state $s$.
Agent-specific counterfactuals allow us to estimate an individual agent’s contribution to NSE. Note that this is a typical form of counterfactual state generation for credit assignment, in that the agent contribution is estimated by either removing it or replacing its actions, while the other agents’ behaviors are fixed [??].
@@
Counterfactual neighbors vs state counterfactuals.
Let ${s}^{i,v}_{c}\subset s^{v}_{c}$ be the set of valid (reachable) counterfactual neighbors for agent $i$. The blame $B_{i}$ for agent $i$ in the joint state $s$ is calculated as,
@@
where $R_{N}$ is the joint NSE penalty function, $R_{N}^{*}=\max_{s\in S}R_{N}(s)$ denotes the maximum joint NSE penalty possible which is used to rescale blame in the range $[0,R_{N}^{*}]$ to avoid impractical values that might be negative or exceed the joint NSE penalty itself. $\epsilon$ is a small fixed value to avoid singularities during normalization and rescaling, and $b_{i}(s)$ is an intermediate value used to calculate $B_{i}(s)$. The above equation assigns blame proportional to the agent’s ability to mitigate NSEs. This ensures that agents already making their best efforts are penalized less compared to other agents.
@@
The agents then solve the MASE-MDP, with the prescribed $R_{1}^{i}$ corresponding to their assigned task and $R^{i}_{N}$ provided by the metareasoner, using lexicographic value iteration (LVI) [?].
Generalizing $R^{i}_{N}$
@@
The penalty function $R^{i}_{N}$ obtained from RECON in Algorithm 1, is calculated based on the blame values corresponding to $\vec{\pi}$ and does not provide any information about potential NSEs that may occur if the agents followed a different joint policy. As a result, NSEs may persist when the agents update their policy by solving the MASE-MDP with $R^{i}_{N}$. To overcome this limitation, we consider a supervised learning approach to generalize the NSE penalty to unseen situations by using the $R^{i}_{N}$, based on the initial $\vec{\pi}$, as training data. The prediction accuracy can further be improved by including the counterfactuals as part of the training [?]. In our experiments, learning to generalize using counterfactuals improves the performance in some settings, compared to generalizing without using counterfactual information. Generalization helps scale the algorithm to systems with large number of agents and eliminates the requirement of having an iterative approach where agents could discover new NSE in unseen states. An iterative RECON will work but it incurs a high computational overhead and does not guarantee a monotonic improvement in policy with each iteration.
@@
(a) Salp agents Refer to caption, inspired from [??], are tasked with collecting physical samples from sites Refer to caption,Refer to caption, and depositing them at lab facility Refer to caption surrounded by corals Refer to caption prone to damage.
@@
(b) Overcooked environment, inspired from [?], shows 4 agents tasked with soup orders. The kitchen shown above has tomatoes Refer to caption, onions Refer to caption, cooking pots Refer to caption, clean dishes Refer to caption, and a serving counter Refer to caption, dirty dishes Refer to caption, and garbage bins Refer to caption for waste.
@@
(c) Warehouse environment, inspired from [?], shows agents Refer to caption tasked with processing shelves Refer to caption, at counter Refer to caption. The narrow corridors Refer to caption with human workers Refer to caption that are inconvenienced around multiple robots carrying shelves.
@@
Figure 2: Example problem instances from domains used in our experiments.
6 Experimental Setup
We evaluate our approach in simulation using three proof-of-concept grid-world domains (salp, overcooked, and warehouse). Figure 2 shows sample instances from each domain. All algorithms are implemented in Python and the simulations were run on an Ubuntu machine with $32$ GB RAM.The MASE-MDP problem is solved using lexicographic value iteration [?] and results are averaged over five instances in each domain. We test with zero NSE tolerance $\eta\!=\!0$ to evaluate the extent of NSE mitigation using RECON, and using $\epsilon=10^{-4}$ for rescaling the blame values.
Baselines
We compare the performance of RECON with four baselines. First is a Naive policy that is optimal for agent task completion and does not optimize NSE mitigation, providing an upper bound on NSE penalty. Second is performing credit assignment using the Difference Reward technique [?] that calculates blame by removing the agent from the system, $B_{i}(s)=R_{N}(s)-R_{N}(s_{-i})$. A direct comparison with this technique is not feasible since it calculates counterfactuals based on assumptions that do not hold in our setting. Therefore, we modify the approach to calculate counterfactuals over dynamic global features of the agents, and calculate blame as $B_{i}(s)=R_{N}(s)-\max_{s^{\prime}\in s^{i,v}_{c}}R_{N}(s^{\prime})$. We use max over all valid counterfactual neighbors to get an aggressive blame assignment, producing conservative (safer) behavior. Third, we compare with a variant of considerate reward inspired from  [?], where they augment agents’ reward functions with value functions of subsequently acting agent to avoid their definition of NSE. Since in our setting, agents operate simultaneously and NSE is defined for multiple agents, we augment agents’ prescribed rewards with blame contributions from the rest of the agents as shown in Equation 4.
$R^{i}_{r}(\hat{s}_{i})=\alpha_{1}\frac{R_{1}^{i}(\hat{s}_{i})}{R_{1}^{*}}+\alpha_{2}\frac{R_{N}(s)-B_{i}(s)}{R_{N}^{*}}$ (4)
where $R_{1}^{*}=\max_{\hat{s}\in\hat{S}_{i}}R_{1}(\hat{s})$ is the maximum possible reward for assigned task for agent $i$, $\alpha_{1}$ and $\alpha_{2}$ are the selfish and care coefficients respectively. $R_{1}^{*}$ and $R_{N}^{*}$ are used to normalization so that their relative scales are not an inherent factor, but a controlled one (using $\alpha_{1}$, $\alpha_{2}$). Fourth, we compare with generalized RECON w/o counterfactual data, which uses only the $R^{i}_{N}$ values from RECON to generalize the penalty situations using supervised learning on global features of the local states and attributed blame values. Finally, we compare with generalized RECON w/ counterfactual (cf) data, which uses both counterfactuals and $R^{i}_{N}$ values to generalize $R^{i}_{N}$ to unseen situations using supervised learning similar to the previous baseline. The techniques are evaluated on the following domains.
Sample Collection using Salps
Salps are marine organisms that can form chains and coordinate movements [?]. We consider a setting where salp-inspired underwater robots must collect chemical samples from different locations in the seabed. Each agent is tasked with collecting a specific type of sample and depositing it to a designated location for analysis. An agent’s state is denoted by $\langle x,y,sample,coral,status\rangle$, where $x,y$ denote agent’s location, $sample$ indicates the sample type, $coral$ indicates presence of coral at $x,y$, and $status$ is a binary variable indicating if the sample has been deposited at the destination. In our experiments, samples can be of type $A,B$ or $X$ indicating no sample. Agents can move in all four directions, and $pick$ and $drop$ samples, each with a reward of $-1$. The reward for task completion is $+100$. The move actions succeed with a probability of $0.8$ or fail and slide to an adjacent location with a probability of $0.2$. Features used for training when generalizing $R^{i}_{N}$ are $\langle sample,coral\rangle$. We test with five instances of $20\!\times\!20$ grids that vary in coral locations. This allows us to evaluate the different approaches to mitigate NSEs, while keeping other conditions fixed.
NSE: The salp-like robots may have chemical residues floating around them when transporting samples. A joint NSE occurs when multiple robots carrying chemical samples are in immediate vicinity of corals, potentially damaging it.
Overcooked
In this domain, agents are tasked with preparing and serving food [?]. The agents work together to deliver a fixed number of tomato and onion soup orders, while keeping the kitchen clean. Each agent is assigned a cleaning or a cooking task, and it receives a reward of $+100$ upon task completion. Actions in non-goal state have a reward of $-1$. In each problem instance, $20\%$ agents are assigned to cleaning and the rest are assigned the cooking task. An agent’s state is represented as $\langle x,y,dir,object,bin,status\rangle$, where $x,y$ denote its location, $dir$ denotes its orientation, $bin$ indicates the presence of garbage bins at $x,y$, and $status$ indicates the completion status of its assigned task. Agents can move forward in all directions, and $interact$ with objects in the kitchen. Interactions vary from picking and putting down ingredients and dishes, to using the cooking pots and dumping garbage in bins. Agent’s move actions succeed with a probability of $0.8$ or slide to an adjacent location with $0.2$ probability. The interact actions succeed with probability of $0.8$ or fail and remain in the same state with probability $0.2$. Features used for generalizing $R^{i}_{N}$ are $\langle object,bin\rangle$. We test with five instances of $15\!\times 15$ grids that vary in the locations of the garbage bins. NSE: The garbage bins emit bad odors and attract flies. Any object involved in soup preparation must therefore be kept away from the garbage bins or the waste must be disposed at a farther bin, otherwise the food quality is affected.
@@ 6 Experimental Setup
Warehouse Inventory Management
[…]
NSE: When multiple robots carrying shelves are simultaneously in the same narrow corridor, it inconveniences human workers to navigate and access the area. The NSE penalty depends on shelf size and the number of agents carrying shelves in the same corridor.
@@
Figure 4: Average NSE penalties and standard deviations, averaged over five problem instances and varying number of agents.
(a) Salp domain
(b) Overcooked domain
(c) Warehouse domain
Figure 5: Average runtime of all techniques, with increasing number of agents across five problem instances, with standard deviation.
NSE Penalty
We employ a logarithmic NSE penalty function (Equation 5) for the domains described earlier. The NSE penalty calculated as,
$R_{N}(s)=\sum_{d\in F_{gd}}\sum_{k\in d}\beta_{k}\cdot\log\left(\alpha_{d}N_{k}+1\right),$ (5)
@@
where $F_{gd}$ is the set of dynamic global features in the environment, $d$ is a feature in $F_{gd}$, and $k$ is the value of feature $d$. $N_{k}$ is the number of agents with dynamic global feature value $d\!=\!k$ in the joint state $s$. $\beta_{k}$ is a scaling factor that is directly related to the NSE penalty for a specific feature value $d=k$. Essentially, higher NSE penalties correspond to higher values of $\beta_{k}$, reflecting that mitigating more critical features yields a more substantial reduction in NSE penalty, aligning with the nuanced dynamics of real-world environments. $\alpha_{d}\,(>0)$ is a sensitivity parameter for NSE associated with feature $d$. A larger $\alpha_{d}$ denotes that the NSE penalty is more sensitive to an increasing number of agents with a specific feature value.
We use a logarithmic function to effectively model scenarios where the NSE impact plateaus with a certain number of agents performing undesirable actions [??]. For example, there may be no significant difference in the NSE impact caused by 10 agents blocking a corridor versus 11 agents blocking a corridor. A log penalty will lead to prioritizing the mitigation of severe NSEs, by updating the policy of a subset of agents. However, note that RECON can work with any penalty function in practice.
7 Results and Discussion
We compare the effectiveness of different techniques in mitigating NSEs, along multiple dimensions such as the number of agents undergoing policy update to mitigate the impact produced, NSE avoidance by generalizing $R^{i}_{N}$ to unseen situations, and run time of the techniques.
Effect of number of agents undergoing policy update     First, we evaluate the effectiveness of our approach in mitigating NSEs, by varying the number of agents whose policy must be updated after $R^{i}_{N}$ is calculated. For each domain, we consider $25$ agents in the environment and vary the percentage of agents undergoing policy update to minimize NSEs, from $10\%$ to $100\%$. We select agents to update policies by ranking them in the decreasing order of their blame values. The results, in Figure 3, show naive policy as a straight line in all the domains, since it is indifferent to NSEs. For the salp sample collection domain, Figure 3(a) shows that the NSE penalty plateaus at the $50\%$ mark, where the difference reward baseline shows its best performance against the naive policy. For the warehouse domain, in Figure 3(c), we see a similar trend of NSE penalty plateauing at the $50\%$ mark however, difference reward fails to mitigate NSE and performs same as the naive policy. In case of the overcooked domain, Figure 3(b) shows a steady decline of NSE penalty using RECON and its generalized versions but similar to the warehouse domain results, difference reward fails to mitigate NSE and performs same as the naive policy.
@@ 7 Results and Discussion
Note that in some cases, NSE may not be avoided even when we update the policies of $100\%$ of the agents in the system. This is because we prioritize completing the task optimally over minimizing NSEs. Hence in some cases, it may be impossible for the agents to avoid NSEs, while optimally completing their tasks. This is a problem characteristic and not a limitation of RECON.
Overall, the results show that RECON is able to mitigate NSEs, without updating the policies of a large number of agents. Based on these results, we update the policies of $50\%$ of agents in all the following experiments, across all domains, since we consistently see a plateau and drop in the NSE penalty at $50\%$. For some settings, updating all the agents’ policies may be effective but it is not practically feasible.
@@
Effect of generalizing $R^{i}_{N}$    Since the penalty function $R^{i}_{N}$ is calculated based on the blame values corresponding to $\vec{\pi}$, policy updates may produce novel NSEs that were not encountered earlier. We evaluate the effectiveness of generalizing $R^{i}_{N}$ to unseen situations by comparing the performances of RECON, with and without generalization using counterfactual data, against the baselines. Based on results from Figure 3, we conduct experiments with $50\%$ of agents undergoing policy update for each technique. Figure 4 shows that both the generalized versions of RECON outperform other methods, consistently reducing NSE penalty across domains. In few instances, generalized RECON with counterfactual data outperforms the generalization without counterfactual data. These results show that generalization is useful and can mitigate NSEs considerably even when counterfactual data is not used for training. We also tested the techniques when $100\%$ of agents undergo policy update. We observe that the NSE penalties further reduce by over $60\%$. While this may be practically infeasible for large systems, it shows the effectiveness of our approach in mitigating NSEs.
Scalability     We evaluate the scalability of our approach in terms of run-time to solve the problems as we increase the number of agents in the system from $10$ to $100$. Figure 5 plots the runtime of different techniques, in the three domains, with varying number of agents. The run times increase in an approximately linear manner as the number of agents increases. The considerate reward baseline takes the least time possible because it involves solving only for one agent as other agents are treated as part of the environment. Difference reward baseline takes longer since policy computation is done in two stages but it does not evaluate all counterfactual neighbors and defaults to the counterfactual producing maximum NSE penalty. RECON takes longer than difference reward since it involves computing and comparing counterfactuals. As expected, RECON with $R_{N}$ generalizations take longer than the other approaches, since they generalize NSE penalties over the entire state space. However, the time increase is small, compared to the reduction in NSE penalties they offer, even in settings with $100$ agents. Since RECON is an offline approach, the slightly higher runtimes for additional processing to generalize $R_{N}$ may be tolerable in many settings. Since RECON is an offline approach, the slightly higher runtimes for additional processing to generalize $R^{i}_{N}$ may be tolerable in many settings.
8 Summary and Future Work

### v2

@@ IV PENALTY DECOMPOSITION
IV-A Blame Estimation Using Counterfactual Evaluation
A blame value is calculated corresponding to each agent’s contribution to the global penalty, using counterfactual information. Since the NSE occurrence is determined only by the dynamic global features $(\vec{f}_{gd})$, counterfactuals must be calculated only over $\vec{f}_{gd}$ to avoid incorrect attribution. For example, consider multiple warehouse robots navigating in a narrow corridor, with some carrying large shelves. The NSE of blocked paths for human access is determined by number of agents carrying large shelves. Generating counterfactuals by considering all state features, including changing agent location, does not provide information on NSE associated with multiple agents carrying large shelves. We therefore estimate blame using counterfactual neighbors, which are a subset of counterfactual states.
Definition 4.
The counterfactual neighbors of a joint state $s$, denoted by $s_{c}$, are the set of all states that differ only in the values of dynamic global features $(\vec{f}_{gd})$ while all other feature values are same as in $s$.
The set of counterfactual neighbors that are reachable from the start state in the environment are referred to as valid counterfactual neighbors and are denoted by ${s}_{c}^{v}$.
Definition 5.
Agent-specific counterfactual neighbors, denoted by $s^{i}_{c}\!\subset s_{c}$, are states that differ in those dynamic global features that can be controlled by agent $i$, while other feature values are same as in the current joint state $s$ (e.g. a warehouse robot evaluated with a different sized box).
Agent-specific counterfactuals estimate an agent’s contribution to NSE, similar to credit assignment where the agent contribution is estimated by either removing it or replacing its actions, while keeping others’ behaviors fixed [15, 17]. Let ${s}^{i,v}_{c}\subset s^{v}_{c}$ be the set of valid (reachable) counterfactual neighbors for agent $i$. The blame $B_{i}$ for agent $i$ in the joint state $s$ is calculated based on the difference between the current NSE penalty and the minimum NSE penalty that could have been achieved by the agent:
@@ IV-A Blame Estimation Using Counterfactual Evaluation
where $R_{N}$ is the joint NSE penalty function, $R_{N}^{*}=\max_{s\in S}R_{N}(s)$ denotes the maximum joint NSE penalty possible which is used to rescale blame in the range $[0,R_{N}^{*}]$ to avoid impractical values that might be negative or exceed the joint NSE penalty itself. $\epsilon$ is a small fixed value to avoid singularities during normalization and rescaling, and $b_{i}(s)$ is an intermediate value used to calculate $B_{i}(s)$. Equation 1 assigns blame proportional to the agent’s ability to mitigate NSEs. This ensures that agents already making their best efforts are penalized less compared to other agents.
@@
The agents then solve the MASE-MDP using lexicographic value iteration (LVI) [27], with the prescribed $R_{1}^{i}$ and $R^{i}_{N}$ provided by the metareasoner.
@@
Generalizing $R^{i}_{N}$ The penalty function $R^{i}_{N}$, calculated using Equation 3, is based on the blame values corresponding to $\vec{\pi}$ and does not provide any information about potential NSEs that may occur if the agents followed a different joint policy. As a result, NSEs may persist when agents update their policies by solving the MASE-MDP with $R^{i}_{N}$. To overcome this limitation, we consider a supervised learning approach to generalize the NSE penalty to unseen situations by using the $R^{i}_{N}$, based on the initial $\vec{\pi}$, as training data. The prediction accuracy can further be improved by including the counterfactuals as part of the training [28]. Generalization helps scale the algorithm to handle many agents, eliminating the need for an iterative RECON, which incurs a high computational overhead.
@@
(a) Salp agents Refer to caption, inspired from [24, 25], are tasked with collecting physical samples from sites Refer to caption,Refer to caption, and depositing them at lab facility Refer to caption surrounded by corals Refer to caption prone to damage.
@@
(b) Overcooked environment, inspired from [5], shows agents cooking and cleaning in kitchen with tomatoes Refer to caption, onions Refer to caption, cooking pots Refer to caption, clean dishes Refer to caption, serving counter Refer to caption, dirty dishes Refer to caption, and waste bins Refer to caption.
@@
(c) Warehouse environment, inspired from [16], shows agents Refer to caption tasked with processing shelves Refer to caption, at counter Refer to caption. The narrow corridors Refer to caption with human workers Refer to caption.
@@
Figure 3: Instances of environments from (a) salp, (b) overcooked, and (c) warehouse domains used in our experiments.
V EXPERIMENTS
We evaluate in simulation and using mobile robots. Code will be made public after paper acceptance. The MASE-MDP problem is solved using LVI [27], with zero NSE tolerance $\eta\!=\!0$ and using $\epsilon=10^{-4}$ for rescaling the blame values.
@@ V EXPERIMENTS
Figure 5: Average NSE penalties and standard deviations, averaged over five problem instances with varying number of agents.
TABLE II: Time taken (in minutes) to solve problems with 100 agents, averaged over five problem instances in each domain.
Domains Considerate Reward Difference Reward RECON Generalized RECON without CF Generalized RECON with CF
Salp $1.1\pm 0.1$ $2.9\pm 0.3$ $4.0\pm 0.4$ $5.5\pm 0.8$ $5.7\pm 0.8$
Overcooked $121.3\pm 3.8$ $174.4\pm 3.4$ $205.1\pm 4.1$ $254.4\pm 3.2$ $288\pm 6.3$
Warehouse $9.1\pm 0.3$ $10.0\pm 0.4$ $12.4\pm 0.2$ $16.1\pm 0.1$ $19.0\pm 0.4$
Baselines We compare the performance of RECON with:
  1. 1.
    Naive Policy that is optimal for the assigned task and does not optimize NSE mitigation, providing an upper bound on NSE penalty;
  2. 2.
    Difference Reward [17] that has been modified to perform blame assignment only considering dynamic global features, to be consistent with our approach, $B_{i}(s)=R_{N}(s)-\max_{s^{\prime}\in s^{i,v}_{c}}R_{N}(s^{\prime})$;
  3. 3.
    Considerate Reward approach [1] that has been modified to support multiple agents acting simultaneously, $R^{i}_{r}(\hat{s}_{i})=\alpha_{1}\frac{R_{1}^{i}(\hat{s}_{i})}{R_{1}^{*}}+\alpha_{2}\frac{R_{N}(s)-B_{i}(s)}{R_{N}^{*}}$ where $R_{1}^{*}=\max_{\hat{s}\in\hat{S}_{i}}R_{1}(\hat{s})$ is the maximum possible reward for assigned task for agent $i$, $\alpha_{1}$ and $\alpha_{2}$ are the selfish and care coefficients respectively. $R_{1}^{*}$ and $R_{N}^{*}$ are used for normalization so that their relative scales are not an inherent factor, but a controlled one (using $\alpha_{1}$, $\alpha_{2}$);
  4. 4.
    Generalized RECON w/o counterfactual data which generalizes the estimated $R^{i}_{N}$ to unseen states, using supervised learning; and
  5. 5.
    Generalized RECON w/ counterfactual (cf) data which generalizes $R^{i}_{N}$ by including the counterfactuals in the training data, using supervised learning.
NSE Penalty Calculation Our experiments use a logarithmic NSE penalty function to model scenarios where the NSE impact plateaus with a certain number of robots involved [8, 31], such as negligible difference in penalty between 10 and 11 robots blocking a corridor. The penalty is calculated as,
$R_{N}(s)=\sum_{d\in F_{gd}}\sum_{k\in d}\beta_{k}\cdot\log\left(\alpha_{d}N_{k}+1\right)$ (4)
@@
where $d$ is a feature in the set of dynamic global features $F_{gd}$, $N_{k}$ is the number of robots with feature value $d\!=\!k$ in the joint state $s$, and $\beta_{k}$ is a scaling factor for penalty based on a specific feature value $d\!=\!k$. Essentially, higher NSE penalties correspond to higher values of $\beta_{k}$, reflecting that addressing key feature values yields a more substantial reduction in NSE penalty. $\alpha_{d}\,(>0)$ is a sensitivity parameter for NSE associated with feature $d$. A larger $\alpha_{d}$ denotes that the NSE penalty is more sensitive to an increasing number of robots with a specific feature value.
V-A Evaluation in simulation
Experiments are conducted using the following three domains in simulation and the results are averaged over five instances in each domain.
Sample Collection using Salps Salp-inspired [24] underwater robots are tasked with collecting chemical samples of type $A$ or $B$ from different locations in the seabed as illustrated in Fig. 3(a). A robot’s state is denoted by $\langle x,y,sample,coral,status\rangle$, where $x,y$ denote robot’s location, $sample$ indicates the sample type with $X$ indicating no sample, $coral$ indicates presence of coral at $x,y$, and $status$ indicates if the sample has been deposited at the destination. Features used to generalize $R^{i}_{N}$ are $\langle sample,coral\rangle$. We test with five $20\!\times\!20$ grids that vary in coral locations. NSE: The salp-like robots may have chemical residues floating around them when transporting samples. A joint NSE occurs when multiple robots carrying chemical samples are in immediate vicinity of corals, potentially damaging it.
Overcooked As shown in Fig. 3(b), robots are tasked with preparing a fixed number of tomato and onion soup orders, while keeping the kitchen clean [30]. In each problem instance, $20\%$ agents are assigned to cleaning and the rest are assigned cooking tasks. An agent’s state is represented as $\langle x,y,dir,object,bin,status\rangle$, where $x,y$ denote its location, $dir$ is its orientation, $bin$ indicates the presence of garbage bins at $x,y$, and task completion status is denoted by $status$. Agents can move forward in all directions, and $interact$ with objects. Interactions include picking up and putting down objects, cooking, and dumping garbage in bins. Features used for generalizing $R^{i}_{N}$ are $\langle object,bin\rangle$. We test with five $15\!\times 15$ grids with varying locations of the garbage bins. NSE: The garbage bins emit bad odors and attract flies. Any object involved in soup preparation must therefore be kept away from the garbage bins or the waste must be disposed at a farther bin.
[…]
V-B Results and Discussion
@@
Number of agents undergoing policy update  For each domain in simulation, we consider $25$ agents in the environment and vary the percentage of agents undergoing policy update to minimize NSEs, from $10\%$ to $100\%$ (Fig. 4). We select agents to update policies by ranking them in the decreasing order of their blame values. Note that in some cases, NSE may not be avoided even when we update the policies of $100\%$ of the agents in the system. This is because we prioritize completing the task optimally over minimizing NSEs, and it may be impossible for the agents to avoid NSEs while optimally completing their tasks. This is a problem characteristic and not a limitation of RECON. Overall, the results show that RECON and its variants with generalizations can successfully mitigate NSEs, without updating the policies of a large number of agents.
@@
NSE mitigation and scalability Fig. 5 shows that RECON and both version of generalized RECON outperform other methods, consistently reducing NSE penalty across domains and with varying number of agents in the system. The results also show that generalization is useful and can mitigate NSEs considerably even when counterfactual data is not used for training. The above results are with $50\%$ of agents undergoing policy update for each technique, based on the trend in Fig. 4. While updating 100% agents reduced the NSE penalty in our experiments, it is practically infeasible to implement it in practice for large systems.
Table II shows the run time (in minutes) of various techniques to solve problems with $100$ agents. The results indicate an approximately linear increase in the run time with increase in the number of agents. The considerate reward baseline takes the least time as it solves a single-agent problem with other agents treated as part of the environment.
V-C Evaluation using mobile robots
We conduct experiments in the salp domain with two Turtlebots to validate real-time usability and effectiveness of our approach in mitigating NSEs, using the map in Fig. 1. The map layout and operation zone of each robot make it inevitable to fully avoid NSEs. Table III shows the percentage of NSE states encountered, with standard deviation. We report results only with Generalized RECON w/ counterfactual data as it consistently performs similar to or better than the other RECON versions (Fig. 5).
TABLE III: Average NSE encounters and standard deviation from our experiments with two mobile robots as shown in Fig. 1.
Approach Average NSE Encounters
Naive $40.33\%\pm 5.42$
Difference Reward $38.86\%\pm 6.73$
Considerate Reward $46.83\%\pm 4.77$
Gen. RECON w/ cf data $10.33\%\pm 5.33$
Considerate Reward with fewer agents is worse than Naive approach, as each agent is less cautious about avoiding NSEs due to reduced joint penalties. Generalized RECON w/ counterfactual data produces the least NSEs, outperforming the baselines.
VI SUMMARY AND FUTURE WORK